OpenStudy (anonymous):

What is the sum of the arithmetic sequence 9, 14, 19, …, if there are 36 terms? 3,402 3,438 3,474 3,510 What is the sum of the arithmetic sequence 152, 138, 124, …, if there are 24 terms? –240 –216 –192 –168 What is the sum of the arithmetic sequence 308, 304, 300, …, if there are 23 terms? 6,049 6,072 6,095 6,118 What is the 9th term of the geometric sequence 3, 9, 27, …? 2,187 6,561

OpenStudy (anonymous):

google it

OpenStudy (nikato):

plz one question per post. scrolling up and down to see ur question will get annoyin

OpenStudy (nikato):

so do u know the common difference of 9,14,19,...

OpenStudy (anonymous):

9,14,19,....... a=5 c.d d=a2-a1=14-9=5 n=36 \[S=\frac{ n }{ 2 }\left\{ 2a+\left( n-1 \right)*d \right\}\] calculate